OpenStudy (anonymous):

a group of 5 people is to be chosen from eight people: alex betty chris declan emma fui greg and hanna. If each of the possible groups are equally likely what is the probability that the chosen group will contain greg or hanna?

OpenStudy (anonymous):

2/8?

OpenStudy (anonymous):

nah that would be correct if it was choosing from the 8, but its using combinations, because theres many different combinations of a group of 5 using 8 people if you know what i mean :) thanks anyway

OpenStudy (anonymous):

um ok sorry :) I'm not really sure :)

OpenStudy (anonymous):

thats okay thanks for trying!

OpenStudy (anonymous):

haha I've never been very good with those types of problems :)

OpenStudy (anonymous):

Is it possible to have greg and hanna both in a single group?

OpenStudy (anonymous):

I think it might be.....

OpenStudy (anonymous):

hmm yeah i would say it would be possible because its asking greg or hanna :)

OpenStudy (anonymous):

\[2\over 8\] as greg and hanna are 2 ppl out of 8 ppl

OpenStudy (anonymous):

Sample spaces= 2/8 favorable outcome= 2/8

OpenStudy (anonymous):

yeah but its not out of 8 people, you have to use combinations because its making a group of 5 and there is 8nCr5 ways of making 5 groups of people D:

OpenStudy (dumbcow):

look at case where neither greg nor hannah is chosen then subtract that from 1 to obtain compliment \[\rightarrow P = 1-\frac{6C5}{8C5}\]

OpenStudy (anonymous):

@dumbcow u r right!

OpenStudy (anonymous):

yep @dumbcow is right.

OpenStudy (anonymous):

yep that is the answer in the back of my book thankyou dumbcow! :)

OpenStudy (dumbcow):

welcome :)